OpenStudy (anonymous):

What is the 20th term in the arithmetic sequence 200, 190, 180 ...

OpenStudy (yttrium):

See you have here: \[a_{n} = 210-10n\] substituting n = 20, well have \[a_{20} = 220 - 10(20)\] \[a_{20} = 220 - 200 = 20\]

OpenStudy (anonymous):

The sequence function is: -10 (n-21) -10(20-21)=-10(-1)=10

OpenStudy (anonymous):

Subtract 10 until you reach the 20th term in the sequence :/ (lot plainer said that those guys ^ lol)